# Effects of Geometric Modelling and Blood Rheology in Patient-Specific Arterial Blood Flow Simulations with Speed-Accuracy Trade-Off Analysis

### v3

@@ 3.3.2 Correlations
[…]
@@
where $C_{12}$, $C_{13}$ and $C_{23}$ are the correlation parameter between levels 1 and 2, levels 1 and 3 and levels 2 and 3, respectively. By definition, $C_{11}$ should be unity and is used in the plots for validation. Thus, correlation parameters provide a sense of a “hotspot” (in WSS or pressure) being observed at a similar location when two different geometry levels are used. If a simplified geometry is able to capture all hotspots of level 1 at their respective locations, the value of the correlation parameter would be close to unity.
@@
Here, we evaluate correlation levels for WSS and pressure. These correlations are computed at specific time instants of the waveform. The correlation coefficient, $C$, is evaluated for the wall shear stress (WSS) magnitude and wall pressure over the entire cardiac cycle for the selected viscosity models, and the results are presented in Figure 11. The correlation trends for both WSS (Figure 11(a)) and wall pressure (Figure 11(b)) are examined. The WSS values of $C_{12}$ remain close to unity for the non-Newtonian models but show large fluctuations for the Newtonian model. In contrast, $C_{13}$ and $C_{23}$ remain further away from unity. This quantitatively supports our preceding discussion. For WSS, Level 2 predictions remain highly correlated with Level 1 (and thus retain accuracy), whereas much of this accuracy is lost in Level 3. However, the Newtonian model is not well behaved for these simulations, showing large fluctuations in correlation values across models. Particularly, cross-correlations $C_{12}$ and $C_{13}$ are large during phases C and D for the Newtonian simulation, suggesting the role of the flow structures appearing in Level 1. This perhaps indicates that the non-Newtonian models of blood should be used for such simulations in geometries that contain a high degree of irregularity. Conversely, accuracy loss is observed to be mild in pressure values.
@@
Figure 11: Correlations of (a) wall shear stress magnitude and (b) wall pressure are plotted with time for the selected viscosity model, i.e, (i) Apostilidis and Beris, (ii) Carreau-Yasuda and (iii) Newtonian. The correlation parameter values are highlighted at phases of the velocity waveform given at t = 0 s (A), t = 0.35 s (B), t = 0.55 s (C), and t = 0.9 s (D).
3.3.3 Time-averaged wall shear stress (TAWSS)
@@
Along with WSS, time-averaged wall shear stress (TAWSS) is a primary risk parameter to monitor the development and progression of atherosclerosis, which can further lead to various types of cardiovascular diseases. Both low and high-wall shear stress promotes pro-inflammatory and anti-inflammatory responses by the endothelial cells [31]. An average WSS over the cardiac cycle is the time-averaged wall shear stress (TAWSS), which further helps identify the high and low WSS zones. Mathematically, TAWSS is defined by the following relation:
[…]
@@
Here, $T$ is the time period of the cardiac cycle. Time-averaged wall shear stress (TAWSS) calculated for the fourth cycle of the simulations is shown in Figure 12 for the three levels of the geometry of the descending aorta and selected viscosity models. In Figure 12 (i),(ii) and (iii) TAWSS contours are plotted for Apostolidis and Beris, Carreau Yasuda, and Newtonian viscosity models, respectively. The trends are qualitatively similar to WSS discussed earlier. In Level 1, the perturbations and asymmetry of the geometry generate a significant amount of zones for high and low TAWSS. Some of these zones are retained in Level 2 with less asymmetry and a reduction in the shear stress magnitudes. A gradual variation of TAWSS is obtained in Level 3 which is a smooth geometry. The trends are similar across the three viscosity models. However, the perturbations in TAWSS are slightly different with the Apostolidis and Beris model, whereas the profiles are similar for the Newtonian and Carreau-Yasuda models. The values are slightly higher with the Apostolidis and Beris model in Level 1. Thus, again, concerning spots of extreme stress, Level 2 captures the essence partially while variations in Level 3 are averaged out.
@@
Figure 12: Time average wall shear stress (TAWSS) shown for the three viscosity models for Levels 1,2 and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
3.3.4 Oscillatory shear index (OSI)
@@
In pulsatile flow, the shear stress near the artery wall fluctuates, making the endothelial cell deliver a pro-inflammatory response. The Oscillatory Shear Index (OSI) introduced by Ku et al. [34] provides a measure of these fluctuations. The OSI can be calculated as follows:
[…]
@@
OSI varies between 0 and 0.5, the latter representing a purely oscillatory response. The high-value region shows a high time-periodic oscillation of the wall shear stress.
@@
Figure 13: Oscillatory shear index (OSI) shown for the three viscosity models for Levels 1,2 and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
@@
Figure 13 plots Oscillatory Shear Index contours for the selected viscosity models. OSI contours depict the zone of shear oscillation and its intensity. In Figure 13, the OSI distributed zones in the Level 1 artery can be seen with hotspots of high and low OSI values, signifying the nature of shear stress oscillations. The region with high OSI values is more prone to a pro-inflammatory response by the endothelial cells, whereas the low value of the OSI zone indicates unidirectional shear stress; therefore, an anti-inflammatory response for the cell is obtained. As the complexity of the descending aorta gradually decreases, the OSI zones are also averaged uniformly as the geometrical features are toned down in the subsequent levels of simplification. Lastly, a uniform OSI distribution is observed in Level 3, as no geometrical disturbances are present. It is clear that the asymmetry of Level 1 is lost in both Levels 2 and 3 due to the geometric simplification.
@@
A fair comparison of the time-averaged hemodynamic indicators across geometric levels is shown in Fig. 14 for TAWSS (Fig. 14 (a)) and OSI (Fig.14 (b)) averaged over the circumference at each axial location. As might be expected, the predictions in Level 3 are a monotonic function, while predictions in Levels 1 and 2 show oscillations due to axial variations in the geometry. For all data shown, the Level 3 predictions represent an average behavior of the oscillatory predictions obtained for Levels 1 and 2. The usefulness of Level 2 is readily visible in these results since, mostly, it closely follows the predictions of Level 1. In the present discussion, the Level 1 predictions are averaged over the circumference and the Level 2 geometry is built by averaging diameters at various axial locations. The consistency of Levels 1 and 2 is visible even more for the non-Newtonian models and shows some differences for the Newtonian fluid, especially for OSI. These results bolster the proposition that a somewhat smoother geometry (like Level 2) will be able to capture the fluctuations in hemodynamic parameters, although the local asymmetry will be lost. As discussed later in Section 3.5, this presents a valid ground for a speed-accuracy trade-off since convergence in the smoother geometries is faster and yet captures the essence of the predictions of Level 1. However, only Level 1 (patient-specific geometry) can be used without any further simplifications in applications where the details of local asymmetry are required.
@@
Figure 14: Comparison of TAWSS and OSI for (i) Newtonian, (ii) Carreau-Yasuda, and (iii) Apostolidis and Beris models. (a) TAWSS: Time-Averaged Wall Shear Stress. (b) OSI: Oscillatory Shear Index for the geometric simplification levels 1-3 of the descending aorta. Results of Level 1 shown here are averaged over the circumference for each axial location.
@@
For TAWSS, Level 2 is able to provide predictions close to those of the circumferential averaged Level 1 data. However, Level 3 is only able to show a monotonic function representing an overall average trend. It lacks the oscillation and fluctuations that are caused due to area fluctuation, characteristic of Levels 1 and 2. It is possible to ask, from the prediction of TAWSS of Level 3 and knowing the area changes, is it possible to recover the prediction of Levels 1 and 2? The following approach is adopted here. Shear stress is given by
@@ 3.4 $Q$ criterion
$Q=\frac{||\bar{\bar{\Omega}}||-||\bar{\bar{\gamma}}||}{2}$ (14)
A positive $Q$ value shows the region with a dominance of rotational flow. Similarly, a negative value denotes a zone where rotational flows are not dominant. In Figure 15a, an iso-surface of the $Q$-criterion is shown for $Q$ = 0.1. Similarly, in Figure 15b, it is shown for $Q$ = -0.1. It is clear that the simplification of the artery leads to a uniform flow distribution. Strong vortex structures are created in Level 1 and progressively disintegrate in Level 2 and Level 3.
@@
In the deceleration phase, the $Q$ function swells out, filling the tube volume. During the diastolic flow is weak, and the $Q$ contour created in Level 3 occupies a smaller volume within the tube. Overall, vortex structure weakens considerably as the geometry is simplified. A similar trend is also seen for the contours of $Q$ = -0.1, as shown in Figure 15b. The nature of the distribution of the positive values of the $Q$ function is associated with wall shear stress, whereas negative $Q$ values indicate that strain dominates over rotation, signifying zones of strong deformation rather than vortex formation.
3.5 Speed Accuracy trade-off
In the preceding sections, we have shown details of the accuracy aspects arising from various simplification levels. As discussed, Level 2 mostly captures the local effects for all parameters, but the local asymmetry visible in the original Level 1 is lost. Level 3 is unable to capture any of the local effects and only shows a “representative average” of results obtained in Level 1. However, for the wall shear stress, simulation data of Level 3 can be reconstructed using scaling principles, which agree well with the detailed calculations of Level 1. Thus, each level has its own utility and is able to provide limited but useful information about the flow field. In the present section, we discuss the reduction in computational time with geometric simplification. In Level 2 geometry, we used 25 individual slices over the length of the artery, each slice being a cylindrical patch. To perform an analysis of the effect of this number, we reconstructed the entire artery by extracting the cross-sectional area of the Level 1 geometry along its length. We then developed multiple equivalent geometries consisting of 25, 20, 15, 10, and 5 slices apart from Level 3, which does not have any slices. Using a grid-independence test, we analysed all reconstructed geometries with Newtonian blood viscosity under steady flow conditions, employing a uniform inlet velocity.
@@ 3.5 Speed Accuracy trade-off
Figure 16: Speed-accuracy analysis is carried out by performing (a) grid independence test for simplified geometries up to 25 slices and comparing their respective friction factors. (b) The maximum number of iterations required to reach the tolerance of 1e-5 in velocity is plotted as a function of the element size. (c) Maximum number of iterations is plotted for the Apostolidis and Beris (A&\&&B) and Newtonian viscosity models for Level 2 (with 25 slices) and Level 3 simplification.
@@
Further, to show the speed increase with geometric simplification, we show the maximum number of iterations required to reach a prescribed convergence criterion as a function of the element size, in Figure 16(b). The data reveal that simplified geometries with 5 to 20 slices require approximately 150 iterations. In contrast, the simplified artery with 25 slices and the geometry with zero slices need around 300 iterations for the smallest element size and 200 iterations for the largest element size. For Level 1 geometry, the number of iterations increases as the element size is decreased, which is consistent with expectations for unstructured meshes over complex three-dimensional geometries. Thus, there is a clear trend of improvement in terms of the number of iterations (and thus, the speed of computations) as we move towards simplified geometric representations. The difference is as large as four times between Level 1 and Level 2 (with 25 slices) at their optimal element sizes.
@@
In Figure 16(c), the Apostolidis and Beris viscosity model is compared with the Newtonian viscosity model with respect to the maximum number of iterations as a function of element size for Level 2 and Level 3 geometries. A significant difference is observed when the viscosity model becomes non-Newtonian. For the converged mesh, nearly 5000 iterations are required for the non-Newtonian viscosity model as compared to the Newtonian. Overall, for meshes with acceptable grid quality, geometric simplification reduces the number of iterations required, nonlinearity of the viscosity model increases the number of iterations, while grid refinement increases the number of iterations.
4 Discussion

### v4

@@ III.3.2 Correlations
$\theta=\text{WSS}_{c}-\overline{\text{WSS${}_{c}$}};\ \text{for wall shear stress.}$ (12)
Here, $\overline{\text{WSS${}_{c}$}}$ is the average over the length of the artery, circumferential average wall shear stress (WSSc) at any location, and $\theta$ is the deviation parameter. The deviation parameter $\theta$ is calculated for the simplified levels of the descending aorta, i.e., $\theta_{1}$$\theta_{2}$, and $\theta_{3}$ for Level 1, Level 2, and Level 3, respectively. Keeping Level 1 as a reference, the correlation $(C)$ between the different geometry levels is defined as:
@@
where $C_{12}$, $C_{13}$, and $C_{23}$ are the correlation parameters between levels 1 and 2, levels 1 and 3, and levels 2 and 3, respectively. By definition, $C_{11}$ should be unity and is used in the plots for validation. Thus, correlation parameters provide a sense of a “hotspot” (in WSS or pressure) being observed at a similar location when two different geometry levels are used. If a simplified geometry can capture all hotspots of level 1 at their respective locations, the value of the correlation parameter would be close to unity.
@@
Here, we evaluate correlation levels for WSS and pressure. These correlations are computed at specific time instants of the waveform. The correlation coefficient, $C$, is evaluated for the wall shear stress (WSS) magnitude and wall pressure over the entire cardiac cycle for the selected viscosity models, and the results are presented in Figure 12. The correlation trends for both WSS (Figure 12(a)) and wall pressure (Figure 12(b)) are examined. The WSS values of $C_{12}$ remain close to unity for the non-Newtonian models but show large fluctuations for the Newtonian model. In contrast, $C_{13}$ and $C_{23}$ remain further away from unity. This quantitatively supports our preceding discussion. For WSS, Level 2 predictions remain highly correlated with Level 1 (and thus retain accuracy), whereas much of this accuracy is lost in Level 3. However, the Newtonian model is not well-behaved for these simulations, exhibiting large fluctuations in correlation values across models. Particularly, the cross-correlations $C_{12}$ and $C_{13}$ are large during phases C and D for the Newtonian simulation, suggesting the role of the flow structures that appear in Level 1. This suggests that non-Newtonian models of blood may be more suitable for simulations in geometries with a high degree of irregularity. Conversely, accuracy loss is observed to be mild in pressure values.
@@
Figure 12: Correlations of (a) wall shear stress magnitude and (b) wall pressure are plotted with time for the selected viscosity model, i.e, (i) Apostilidis and Beris, (ii) Carreau-Yasuda, and (iii) Newtonian. The correlation parameter values are highlighted at phases of the velocity waveform given at t = 0 s (A), t = 0.35 s (B), t = 0.55 s (C), and t = 0.9 s (D).
III.3.3 Time-averaged wall shear stress (TAWSS)
@@
Along with WSS, time-averaged wall shear stress (TAWSS) is a primary risk parameter to monitor the development and progression of atherosclerosis, which can further lead to various types of cardiovascular diseases. Both low- and high-wall shear stress promote pro-inflammatory and anti-inflammatory responses in endothelial cells [31]. An average WSS over the cardiac cycle is the time-averaged wall shear stress (TAWSS), which further helps identify the high and low WSS zones. Mathematically, TAWSS is defined by the following relation:
$\text{TAWSS}=\frac{1}{T}\int^{T}_{0}|\bar{\tau}_{w}|dt$ (14)
@@
Here, $T$ is the time period of the cardiac cycle. The time-averaged wall shear stress (TAWSS) calculated for the fourth cycle of the simulations is shown in Figure 13 for the three levels of the geometry of the descending aorta and the selected viscosity models. In Figures 13 (i),(ii), and (iii), TAWSS contours are plotted for Apostolidis and Beris, Carreau Yasuda, and Newtonian viscosity models, respectively. The trends are qualitatively similar to those discussed earlier in WSS. In Level 1, the perturbations and asymmetry of the geometry generate a significant number of zones for high and low TAWSS. Some of these zones are retained in Level 2 with less asymmetry and a reduction in the shear stress magnitudes. A gradual variation of TAWSS is observed in Level 3, characterized by a smooth geometry. The trends are similar across the three viscosity models. However, the perturbations in TAWSS are slightly different from those in the Apostolidis and Beris model, whereas the profiles are identical for the Newtonian and Carreau-Yasuda models. The values are somewhat higher with the Apostolidis and Beris model in Level 1. Thus, again, concerning spots of extreme stress, Level 2 captures the essence partially, while variations in Level 3 are averaged out.
@@
Figure 13: Time average wall shear stress (TAWSS) shown for the three viscosity models for the Levels 1,2, and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
III.3.4 Oscillatory shear index (OSI)
@@
In pulsatile flow, the shear stress near the artery wall fluctuates, making the endothelial cell deliver a pro-inflammatory response. The Oscillatory Shear Index (OSI) introduced by Ku et al. [ku1983pulsatile] provides a measure of these fluctuations. The OSI can be calculated as follows:
$\text{OSI}=\frac{1}{2}\left(1-\frac{|\int^{T}_{0}\bar{\tau}_{w}dt|}{\int^{T}_{0}|\bar{\tau}_{w}|dt}\right)$ (15)
@@
OSI varies between 0 and 0.5, the latter representing a purely oscillatory response. The high-value region exhibits a pronounced time-periodic oscillation in the wall shear stress.
@@
Figure 14: Oscillatory shear index (OSI) shown for the three viscosity models for Levels 1,2, and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
@@
Figure 14 plots Oscillatory Shear Index contours for the selected viscosity models. OSI contours depict the zone of shear oscillation and its intensity. In Figure 14, the OSI distributed zones in the Level 1 artery can be seen with hotspots of high and low OSI values, signifying the nature of shear stress oscillations. The region with high OSI values is more prone to a pro-inflammatory response by the endothelial cells, whereas the low OSI value indicates unidirectional shear stress; therefore, an anti-inflammatory response is obtained for the cell. As the complexity of the descending aorta gradually decreases, the OSI zones are also averaged uniformly, as the geometrical features are refined in the subsequent levels of simplification. Lastly, a uniform OSI distribution is observed in Level 3, as no geometrical disturbances are present. The asymmetry of Level 1 is lost in both Levels 2 and 3 due to the simplification of the geometric structure.
@@
A fair comparison of the time-averaged hemodynamic indicators across geometric levels is shown in Fig. 15 for TAWSS (Fig. 15 (a)) and OSI (Fig.15 (b)) averaged over the circumference at each axial location. As might be expected, the predictions in Level 3 are a monotonic function, while predictions in Levels 1 and 2 show oscillations due to axial variations in the geometry. For all data shown, the Level 3 predictions represent an average behavior of the oscillatory predictions obtained for Levels 1 and 2. The usefulness of Level 2 is readily apparent in these results, as it mostly closely follows the predictions of Level 1. In the present discussion, the Level 1 predictions are averaged over the circumference, and the Level 2 geometry is built by averaging diameters at various axial locations. The consistency of Levels 1 and 2 is more pronounced for the non-Newtonian models, revealing some differences for the Newtonian fluid, particularly for OSI. These results support the proposition that a somewhat smoother geometry (like Level 2) can capture fluctuations in hemodynamic parameters, albeit at the expense of local asymmetry. As discussed later in Section 3.5, this presents a valid ground for a speed-accuracy trade-off since convergence in the smoother geometries is faster and yet captures the essence of the predictions of Level 1. However, only Level 1 (patient-specific geometry) can be used without any further simplifications in applications where the details of local asymmetry are required.
@@
Figure 15: Comparison of TAWSS and OSI for (i) Newtonian, (ii) Carreau-Yasuda, and (iii) Apostolidis and Beris models. (a) TAWSS: Time-Averaged Wall Shear Stress. (b) OSI: Oscillatory Shear Index for the geometric simplification levels 1-3 of the descending aorta. The results of Level 1 shown here are averaged over the circumference for each axial location.
@@
For TAWSS, Level 2 can provide predictions that are close to those of the circumferentially averaged Level 1 data. However, Level 3 is only able to display a monotonic function that represents an overall average trend. It lacks the oscillation and fluctuations caused by area fluctuations, characteristic of Levels 1 and 2. It is possible to ask, from the prediction of TAWSS of Level 3 and knowing the area changes, is it possible to recover the prediction of Levels 1 and 2? The following approach is adopted here. Shear stress is given by
@@ III.4 $Q$ criterion
[…]
A positive $Q$ value indicates a region dominated by rotational flow. Similarly, a negative value denotes a zone where rotational flows are not dominant. In Figure 16a, an iso-surface of the $Q$-criterion is shown for $Q$ = 0.1. Similarly, in Figure 16b, it is shown for $Q$ = -0.1. The simplification of the artery leads to a uniform flow distribution. Strong vortex structures are created in Level 1 and progressively disintegrate in Level 2 and Level 3.
@@
In the deceleration phase, the $Q$ function swells out, filling the tube volume. During diastolic flow, the $Q$ contour created in Level 3 occupies a smaller volume within the tube. Overall, the vortex structure weakens considerably as the geometry is simplified. A similar trend is also seen for the contours of $Q$ = -0.1, as shown in Figure 16b. The nature of the distribution of the positive values of the $Q$ function is associated with wall shear stress. In contrast, negative $Q$ values indicate that strain dominates over rotation, signifying zones of intense deformation rather than vortex formation.
III.5 Speed Accuracy trade-off
In the preceding sections, we have shown details of the accuracy aspects arising from various simplification levels. As discussed, Level 2 mostly captures the local effects for all parameters, but the local asymmetry visible in the original Level 1 is lost. Level 3 is unable to capture any of the regional effects and only shows a “representative average” of the results obtained in Level 1. However, for the wall shear stress, simulation data of Level 3 can be reconstructed using scaling principles, which agree well with the detailed calculations of Level 1. Thus, each level has its own utility and can provide limited but valuable information about the flow field. In this section, we discuss the reduction in computational time achieved through geometric simplification. In Level 2 geometry, we used 25 individual slices over the length of the artery, each slice being a cylindrical patch. To analyze the effect of this number, we reconstructed the entire artery by extracting the cross-sectional area of the Level 1 geometry along its length. We then developed multiple equivalent geometries consisting of 25, 20, 15, 10, and 5 slices, apart from Level 3, which has no slices. Using a grid-independence test, we analysed all reconstructed geometries with Newtonian blood viscosity under steady flow conditions, employing a uniform inlet velocity.
@@ III.5 Speed Accuracy trade-off
Figure 17: Speed-accuracy analysis is carried out by performing (a) grid independence test for simplified geometries up to 25 slices and comparing their respective friction factors. (b) The maximum number of iterations required to reach the tolerance of 1e-5 in velocity is plotted as a function of the element size. (c) The maximum number of iterations is plotted for the Apostolidis and Beris (A&\&B) and Newtonian viscosity models for Level 2 (with 25 slices) and Level 3 simplification.
@@
Further, to show the speed increase with geometric simplification, we show the maximum number of iterations required to reach a prescribed convergence criterion as a function of the element size, in Figure 17(b). The data reveal that simplified geometries with 5 to 20 slices require approximately 150 iterations. In contrast, the simplified artery with 25 slices and the geometry with zero slices need around 300 iterations for the smallest element size and 200 iterations for the largest element size. For Level 1 geometry, the number of iterations increases as the element size is decreased, which is consistent with expectations for unstructured meshes over complex three-dimensional geometries. Thus, there is a clear trend of improvement in terms of the number of iterations (and, consequently, the speed of computation) as we move towards simplified geometric representations. The difference is as large as four times between Level 1 and Level 2 (with 25 slices) at their optimal element sizes.
@@
In Figure 17(c), the Apostolidis and Beris viscosity model is compared with the Newtonian viscosity model with respect to the maximum number of iterations as a function of element size for Level 2 and Level 3 geometries. A significant difference is observed when the viscosity model becomes non-Newtonian. For the converged mesh, nearly 5,000 iterations are required for the non-Newtonian viscosity model compared to the Newtonian model. Overall, for meshes with acceptable grid quality, geometric simplification reduces the number of iterations required, while the nonlinearity of the viscosity model increases the number of iterations; grid refinement, on the other hand, increases the number of iterations.
IV Discussion